### v1

# Maximum entropy principle for quantum processes

Abstract
The maximum entropy principle states that the maximum entropy among all quantum states with a fixed mean energy is achieved only by the thermal state of given mean energy. In this notes, we prove the maximum entropy principle for quantum processes— the entropy of a quantum channel with fixed mean energy is maximum if and only if the channel is absolutely thermalizing channel with the fixed output thermal state of that mean energy. This allows for an alternate approach to describe emergence of the absolute thermalization processes under energy constraints in the observable universe.
I Introduction
Entropy is a primitive information-theoretic quantity that measures randomness or uncertainty in the system. Entropy of a quantum process quantifies the minimum uncertainty in the output of the process conditioned on the reference system Devetak et al. (2006); Gour and Wilde (2021). It has implications on the erasure cost of the logical output of a bipartite unitary circuit when the eraser has access to the ancillary output Badhani et al. (2025). In quantum information theory, quantum channels describe quantum processes that are completely positive, trace-preserving linear maps. The basic randomness and information measures like entropy, conditional entropy, mutual information, and conditional mutual information of quantum channels provide insight to their underlying causal structures and characteristics Das et al. (2024); Badhani et al. (2025). In this short notes, we state and prove the maximum entropy principle for quantum channels which can be deemed as channelized version of the well-known maximum entropy principle for quantum states Jaynes (1980):
Theorem (Informal statement for Theorem 1).
The maximum entropy principle for quantum processes states that among all channels with a fixed mean output energy $E$, i.e., the maximum output mean energy of a channel is $E$, the channel entropy is maximum if and only if the channel is absolutely thermalizing channel with a fixed output thermal state with mean energy $E$.
See Fact 1 and Theorem 1 for the formal statements of the principle for the quantum states and channels, respectively.
I.1 Preliminaries
We first set the standard notations, recall standard definitions, and known facts from the literature.
We consider separable Hilbert spaces. Let $\operatorname{St}(A)$ and $\operatorname{Pos}(A)$ denote the sets of all density operators and positive semidefinite operators, respectively, defined on the Hilbert space $\mathcal{H}_{A}$. The Hilbert space $\mathcal{H}_{AB}$ of a composite system $AB$ is given as $\mathcal{H}_{AB}:=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. We may simply denote $\mathcal{H}_{A}$ as $A$. The dimension of $A$ is $|A|:=\dim(A)$. $\mathbbm{1}_{A}$ is the identity operator on $A$. $\omega_{A}$ denotes operator on $A$ and for brevity we may denote it as $\omega$. For $\rho\in\operatorname{St}(AB)$, $\rho_{B}=\operatorname{tr}_{A}(\rho_{AB})$ is the reduced state on $B$. A quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is a completely positive, trace-preserving linear map with input on $A^{\prime}$ and output on $A$. For a linear map $\mathcal{N}_{A^{\prime}\to A}$ and $\psi\in\operatorname{St}(RA^{\prime})$, $\mathcal{N}(\psi_{RA^{\prime}})=\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})$. Let $\operatorname{Ch}(A^{\prime},A)$ denote the set of all quantum channels $\mathcal{N}_{A^{\prime}\to A}$. Let $\mathcal{R}_{A^{\prime}\to A}$ denote the uniformly mixing map, $\mathcal{R}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\operatorname{tr}(\rho_{A^{\prime}})\mathbbm{1}_{A}$ for all $\rho\in\operatorname{Pos}(A^{\prime})$. A replacer channel $\mathcal{R}^{\omega}_{A\to A^{\prime}}$ always outputs a fixed state $\omega_{A}$ for all possible input states $\rho_{A^{\prime}}$, i.e., , $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\omega_{A}$ for all $\rho\in\operatorname{St}(A^{\prime})$.
The Choi operator of a linear map $\mathcal{N}_{A^{\prime}\to A}$ is $\Gamma^{\mathcal{N}}_{RA}=\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\Gamma_{RA^{\prime}})$, where $R\simeq A^{\prime}$ and $\operatorname{id}_{R}$ is the identity supermap with input-output $R$ and $\Gamma_{RA^{\prime}}:=\sum_{i,j=0}^{d-1}\ket{ii}\bra{jj}_{RA^{\prime}}$ with $d=\min\{|R|,|A^{\prime}|\}$. For $d=\min\{|R|,|A|\}<+\infty$, $\Phi_{RA}:=\frac{1}{d}\Gamma_{RA}$ is a maximally entangled state and $\Phi^{\mathcal{N}}_{RA}:=\frac{1}{|A|}\Gamma^{\mathcal{N}}_{RA}$ is a state (Choi state of $\mathcal{N}_{A^{\prime}\to A}$) if $\mathcal{N}_{A^{\prime}\to A}$ is a quantum channel. The Choi operator of a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$ is a product operator $\mathbbm{1}_{R}\otimes\omega_{A}$.
The von Neumann entropy, simply referred to as entropy, of a quantum state $\rho_{A}$ is $S(A)_{\rho}:=-\operatorname{tr}[\rho\log\rho]$. The quantum relative entropy between $\rho\in\operatorname{St}(A),\sigma\in\operatorname{Pos}(A)$ is given by $D(\rho\|\sigma):=\operatorname{tr}\left[\rho(\log\rho-\log\sigma)\right]$ if $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$ else it is +$\infty$+ ∞. $D(\rho\|\sigma)=0$ if and only if $\rho=\sigma$. We note that $S(A)_{\rho}=-D(\rho_{A}\|\mathbbm{1}_{A})$. For a quantum state $\rho\in\operatorname{St}(AB)$, the von Neumann conditional entropy $S(A|B)_{\rho}$ quantifies uncertainty of $A$ conditioned on $B$, $S(A|B)_{\rho}=-\inf_{\sigma\in\operatorname{St}(B)}D(\rho_{AB}\|\mathbbm{1}_{A}\otimes\sigma_{B})$, and $S(A|B)_{\rho}=S(AB)_{\rho}-S(B)_{\rho}$ Cerf and Adami (1997).
Let $\widehat{H}_{A}$ denote the Hamiltonian of quantum system $A$. Throughout this notes, we consider the Hamiltonian $\widehat{H}_{A}$ to be bounded. The mean energy of $\rho\in\operatorname{St}(A)$ is given by $\langle\widehat{H}\rangle_{\rho}:=\operatorname{tr}[\widehat{H}\rho]$. For a Hamiltonian $\widehat{H}_{A}$, the thermal state $\gamma^{\beta}$ is given as $\gamma^{\beta}_{A}={\exp(-\beta\widehat{H}_{A})}/{Z^{\beta}_{A}}$ Lenard (1978), where $\beta=(k_{B}T)^{-1}$ is inverse temperature and $Z^{\beta}_{A}:=\operatorname{tr}[\exp(-\beta\widehat{H}_{A})]$. The temperature $T$ (or, $\beta$) is uniquely determined by the mean energy $\langle\widehat{H}\rangle_{\gamma^{\beta}}$ of the thermal state $\gamma^{\beta}$.
Fact 1 (Jaynes (1980)).
Consider a quantum system $A$ with Hamiltonian $\widehat{H}_{A}$. The maximum entropy principle states that
$\max_{\rho\in\operatorname{St}(A):~{}\operatorname{tr}[\widehat{H}\rho]=E}S(\rho)=S(\gamma^{\beta})$ (1)
for
$\gamma^{\beta}:=\frac{\exp(-\beta\widehat{H})}{\operatorname{tr}[\exp(-\beta\widehat{H})]},\quad\langle\widehat{H}\rangle_{\gamma^{\beta}}=E.$ (2)
Let $\mathcal{R}^{\mathbbm{1}}_{A^{\prime}\to A}$ denote the uniformly mixing map, $\mathcal{R}^{\mathbbm{1}}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\operatorname{tr}(\rho_{A^{\prime}})\mathbbm{1}_{A}$ for all $\rho\in\operatorname{Pos}(A^{\prime})$. The absolutely thermalizing channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ is a replacer channel $\mathcal{R}^{\gamma^{\beta}}_{A^{\prime}\to A}$ that outputs a thermal state $\gamma^{\beta}_{A}$ for all possible input states.
Entropy of a quantum channel.— The quantum relative entropy between two quantum channels $\mathcal{N}_{A^{\prime}\to A},\mathcal{M}_{A^{\prime}\to A}$ is defined as $D[\mathcal{N}\|\mathcal{M}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{M}(\psi_{RA^{\prime}}))$ Leditzky et al. (2018), and it suffices to consider optimization over pure states $\psi_{RA^{\prime}}$ with $R\simeq A^{\prime}$. The (von Neumann) entropy of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as $S[\mathcal{N}]:=-D[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]=\inf_{{\psi}\in\operatorname{St}(RA^{\prime})}S(A|R)_{\mathcal{N}(\psi)}$ Gour and Wilde (2021)(see also Devetak et al. (2006); Yuan (2019), in particular (Sohail et al., 2025, Definition 6)), suffices to optimize over pure $\psi_{RA^{\prime}}$ states. For an arbitrary replacer channel $\mathcal{R}^{\omega}$, its minimum output entropy $\inf_{\rho\in\operatorname{St}(A^{\prime})}S(\mathcal{N}(\rho))$ Giovannetti et al. (2004); De Palma et al. (2017) is equal to its entropy,
$S[\mathcal{R}^{\omega}]=S(\omega)=\inf_{\rho\in\operatorname{St}(A^{\prime})}S(\mathcal{N}(\rho)).$ (3)
II Main Result
We now give a formal definition of the mean energy of a quantum channel and then prove the maximum entropy principle for quantum channels.
Definition 1.
The mean energy of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as the maximum mean energy of the output of the given channel. Let $\widehat{H}_{A}$ be the bounded Hamiltonian of $A$, then
$\langle\widehat{H}\rangle_{\mathcal{N}}:=\sup_{\rho\in\operatorname{St}(A^{\prime})}\operatorname{tr}[\widehat{H}_{A}\mathcal{N}(\rho_{A^{\prime}})].$ (4)
For a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, we have
$S[\mathcal{R}^{\omega}]=S(\omega),\quad\langle\widehat{H}\rangle_{\mathcal{R}^{\omega}}=\langle\widehat{H}\rangle_{{\omega}}.$ (5)
We now state and prove the maximum entropy principle for quantum processes.
Theorem 1 (Maximum entropy principle for quantum processes).
The maximum entropy among all quantum channels $\mathcal{N}_{A^{\prime}\to A}$ with a fixed mean energy $\langle\widehat{H}\rangle_{\mathcal{N}}=E$, is attained if and only if the channel is absolutely thermalizing channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ with $\langle\widehat{H}\rangle_{\gamma^{\beta}}=E$,
$\max_{\begin{subarray}{c}\ \mathcal{N}\in\operatorname{Ch}(A^{\prime},A):\\ \ \langle\widehat{H}\rangle_{\mathcal{N}}=E\end{subarray}}S[\mathcal{N}]=S[\mathcal{T}^{\beta}]=S(\gamma^{\beta}).$ (6)
Proof.
For a quantum state $\rho_{AB}$, we know that $S(A)_{\rho}\geq S(A|B)_{\rho}$; the inequality is saturated $S(A)_{\rho}=S(A|B)_{\rho}$ if and only if $\rho_{AB}$ is a product state $\rho_{A}\otimes\rho_{B}$. This is because $S(A)_{\rho}=S(A|B)_{\rho}$ if and only if $D(\rho_{AB}\|\rho_{A}\otimes\rho_{B})=0$. For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, we have Gour and Wilde (2021); Devetak et al. (2006)
$S[\mathcal{N}]$ $=\inf_{\rho\in\operatorname{St}(RA^{\prime})}\left[S(RA)_{\mathcal{N}(\rho)}-S(R)_{\rho}\right]$ (7)
$\leq\inf_{\rho\in\operatorname{St}(A^{\prime})}S(A)_{\mathcal{N}(\rho)}.$ (8)
It follows from Ineq. (8) that for an arbitrary channel $\mathcal{N}_{A^{\prime}\to A}$, $S[\mathcal{N}]\leq S(\mathcal{N}(\rho_{A^{\prime}}))$ for all $\rho\in\operatorname{St}{(A^{\prime})}$. If $\rho^{(1)}\in\operatorname{St}{(A^{\prime})}$ such that $\langle\widehat{H}\rangle_{\mathcal{N}}=\operatorname{tr}[\widehat{H}_{A}\mathcal{N}(\rho^{(1)}_{A^{\prime}})]$ then also $S[\mathcal{N}]\leq S(\mathcal{N}(\rho^{(1)}_{A^{\prime}}))$. Whereas, for a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, $S[\mathcal{R}^{\omega}]=S(\mathcal{N}(\rho_{A^{\prime}}))=S(\omega)$ for all $\rho\in\operatorname{St}{(A^{\prime})}$.
From Eq. (7) and Ineq. (8), we notice that $S[\mathcal{N}]=\inf_{\rho\in\operatorname{St}(A^{\prime})}S(A)_{\mathcal{N}(\rho)}$ if and only if the channel decorrelates $A$ from $R$, meaning $A$ and $R$ are in product state, for all possible $\rho\in\operatorname{St}{(RA^{\prime})}$. This happens only when the channel $\mathcal{N}$ is a replacer channel, i.e., $S[\mathcal{N}]=\inf_{\rho\in\operatorname{St}(A^{\prime})}S(A)_{\mathcal{N}(\rho)}$ if and only if $\mathcal{N}$ is a replacer channel.
We now note that for every quantum channel $\mathcal{N}_{A^{\prime}\to A}$ with $\langle\widehat{H}\rangle_{\mathcal{N}}=E$, there exists some replacer channels $\mathcal{R}^{\omega(E)}_{A^{\prime}\to A}$, where $\omega(E)$ is to denote $\langle\widehat{H}\rangle_{\omega}=E$. Also, for every quantum channel $\mathcal{N}_{A^{\prime}\to A}$ with $\langle\widehat{H}\rangle_{\mathcal{N}}=E$, there exists an absolutely thermalizing channel $\mathcal{T}^{\beta(E)}_{A^{\prime}\to A}$, where $\beta(E)$ is to denote $\langle\widehat{H}\rangle_{\gamma^{\beta}}=E$.
It directly follows from Fact 1 and $S[\mathcal{R}^{\omega}]=\omega$ that among all replacer channels $\mathcal{R}^{\omega}_{A^{\prime}\to A}$ with $\langle\widehat{H}\rangle_{\mathcal{R}^{\omega}}=E$, the entropy is maximum if and only if the replacer channel is absolutely thermalizing $\mathcal{T}^{\beta}$ and $\langle\widehat{H}\rangle_{\gamma^{\beta}}=E$. This concludes the proof. ∎
The theorem statement above can be generalized to the cases where chemical potentials Haag et al. (1967); Pathria and Beale (2021) are also considered. Fact 1 will get modified, and Theorem 1 will be revised accordingly. We leave formal statement and rigorous proof for future work. We believe our work provides an axiomatic approach based on the maximum entropy principle to describe absolute thermalization and will be of interest to study quantum to classical transitions Popescu et al. (2006); Leverrier and Cerf (2009); Riera et al. (2012); Brandao and Harrow (2013); Mark et al. (2024).
Acknowledgements.
The authors thank Uttam Singh and Himanshu Badhani for discussions. S.D. acknowledges support from the Science and Engineering Research Board (now ANRF), Department of Science and Technology, Government of India, under Grant No. SRG/2023/000217 and the Ministry of Electronics and Information Technology (MeitY), Government of India, under Grant No. 4(3)/2024-ITEA. S.D. also acknowledges support from the National Science Centre, Poland, grant Opus 25, 2023/49/B/ST2/02468. S.D. thanks Harish-Chandra Research Institute, Prayagraj (Allahabad), India and University of Gdańsk, Gdańsk, Poland for the hospitality during his visits.